Project

General

Profile

Bug #1878

Readiness of trench cooling water circuit

Added by Nebrensky, Henry almost 6 years ago. Updated over 5 years ago.

Status:
Closed
Priority:
Normal
Category:
Beam Line
Start date:
26 September 2016
Due date:
% Done:

100%

Estimated time:

Description

Tasks remaining to complete the trench cooling water circuit:


Files

Grundfos_Manual_CRE_CRIE_CRNE_CRTE_SPKE_MTRE_CME_BMS_-_1662523.pdf (5.65 MB) Grundfos_Manual_CRE_CRIE_CRNE_CRTE_SPKE_MTRE_CME_BMS_-_1662523.pdf Trench water - pump manual Nebrensky, Henry, 21 November 2016 12:08
204-72021-2.pdf (122 KB) 204-72021-2.pdf Nebrensky, Henry, 21 November 2016 13:09
204-72021-3.pdf (85.3 KB) 204-72021-3.pdf Nebrensky, Henry, 21 November 2016 13:09
204-72021-1a.pdf (126 KB) 204-72021-1a.pdf Nebrensky, Henry, 21 November 2016 18:07

Related issues

Related to Operations - Bug #1875: Conventional magnet cooling débâcleClosedNebrensky, Henry21 September 2016

Actions
#1

Updated by Nebrensky, Henry almost 6 years ago

To-do list for TRENCH circuit:

  • Conductivity meter
  • PLC (auto vs. manual)
    • Temperature limit and trip in PLC
    • Readout to MLCR
  • Mechanical support for the filter
  • Extend trench water leak sensor needs readout #1889
  • Archiving and alarms: these now go via the Beamline spreadsheet #1880 as the Loading-Bay chiller/de-min loop are for the conventional magnets only
#4

Updated by Nebrensky, Henry over 5 years ago

The pumps are fitted with 0-16 bar sensors, but they cannot provide much more than 7 bar as the flow is too high.

Pump 1 was showing an alarm light. Paul Masterson connected to the pump via the IR dongle and found the alarm was due to the input signal: "Setpoint Signal Out Of Range", most recently at 22:38 on 19 November 2016, and before that 04:38 on 19 November 2016. We were able to re-set the error (eLog 4219 ). These times correspond to particularly low drops in supplied pressure. We hypothesize that the 4-20 mA setpoint connection is poor and occasionally drops out completely (0 mA, which is indeed out of range).

Disconnecting the 4-20 mA connection from the pump caused it to drop down into a low-flow idle state. We had to locally configure the pumps themselves to ignore external set values, after which it was possible to configure them with a local set value via the IR dongle. We have demonstrated that this value persists even if the pump is power cycled. It will be lost if the arrow buttons on the pump are used

We set a pressure of about 5.9 bar within each pump, giving a circuit pressure of 5.3 bar. We then demonstrated that if the pump was powered off and then powered back on it would start up and run at the appropriate pressure, for both pumps. This means that we are now able to swap easily between the two pumps, just by powering only the desired pump.

#5

Updated by Nebrensky, Henry over 5 years ago

Craig has ordered an R100 remote control unit - this means that MICE should be able to query the pumps for their status and alarm history ourselves, rather than needing to wait for an expert.

#6

Updated by Nebrensky, Henry over 5 years ago

Yesterday we increased the flows through Q5 and Q8 which caused the circuit pressure to fall to 5.2 bar; then used R100 to increase pump 2 setpoint to 6 bar to get circuit back to 5.4 bar.

Have also found "Setpoint Signal Out Of Range" errors in pump 2 logs.

#7

Updated by Nebrensky, Henry over 5 years ago

See eLog thread from entry 4389

Essentially, seem to have gradually been losing PLC/C&M functionality, to the extent that we now have one pump completely separated from the control system so that we can control it directly, and no working monitoring of the circuit.

#8

Updated by Nebrensky, Henry over 5 years ago

On 2nd Feb we asked Rob Searle to come down and help Ajit understand the communications between the soft-IOC and the Trench Water
PLC as we were only seeing zeroes in the EPICS panel. The issue turned out to be that some of the data within the
controller had somehow been overwritten with zeroes, though the program itself had remained (through several power
cycles), and the PLC was therefore returning zeroes to the IOC.
Rob Searle reset the PLC and we got the IOC and control panel working. ( eLog 4525 )

Since then the PLC appears to have been working, or at least we've been getting the feedback (pump pressure, etc.) and no problems from the Trench water side during the data run.

#9

Updated by Nebrensky, Henry over 5 years ago

  • Status changed from Open to Closed
  • % Done changed from 0 to 100

All seemed to be working until we started running - then the mixing valve seemed to flip polarity again - 4971

This was NOT associated with any IOC restart or such.

It looks as though the PLC gradually "degrades" for some reason - probably this should be in a new ticket. The rest of the system is fine.

Also available in: Atom PDF